# Interpretability by Design for Efficient Multi-Objective Reinforcement Learning

### v1

@@ 4.5 Ablation Study
The LLE-MORL integrates a locally linear extension process with a subsequent fine-tuning stage. To understand the distinct contributions of these components to the overall performance, our ablation study separates them. We first evaluate LLE-MORL-0, which solely employs the extension process without fine-tuning. As detailed in Table 1 and Table 2, LLE-MORL-0 itself demonstrates competitiveness, achieving strong Hypervolume (HV) and Expected Utility (EU) scores that are often competitive with or superior to baselines. This emphasises the efficacy of our extension mechanism in rapidly discovering a high-quality approximation of the Pareto front.
Subsequently, we assess the improvement of the fine-tuning stage by comparing LLE-MORL (which includes fine-tuning) to LLE-MORL-0. This comparison reveals that the inclusion of fine-tuning consistently yields further improvements in Hypervolume (HV) and Expected Utility (EU) across both sample-efficient and standard-training settings. The impact on Sparsity (SP) is less uniform, which is an expected outcome, as refining solutions towards a more optimal Pareto front can alter their relative spacing. Nevertheless, the consistent enhancements in HV and EU prove the value of fine-tuning for improving the overall quality of the approximate Pareto front and its coverage by diverse solutions. This demonstrates that the extension process provides a strong foundation for the fine-tuning stage that enables LLE-MORL to outperform other algorithms.
@@ 5 Related work
Prior work in Multi-Objective Reinforcement Learning (MORL) offers various strategies for handling conflicting objectives. These can be broadly grouped into single-policy methods and multi-policy methods for approximating the Pareto front. Single-policy approaches, a foundational strategy in MORL, typically convert the multi-objective problem into a single-objective task using a predefined preference or weighting scheme to find a policy optimal for that specific trade-off. A common instance of such a weighting scheme is linear scalarization  (Van Moffaert et al., 2013). Limitations of linear scalarization, particularly in capturing non-convex Pareto fronts, have been addressed by more advanced scalarization functions such as Chebyshev methods (Van Moffaert et al., 2013) and hypervolume-based approaches (Zhang and Golovin, 2020). Further theoretical work has aimed at enhancing scalarization robustness and performance, for instance, by proposing the addition of concave terms to rewards (Lu et al., 2023). Concurrently, significant efforts have developed generalised single-policy models conditioned on preference inputs to achieve adaptability across diverse objectives (Teh et al., 2017; Yang et al., 2019; Basaklar et al., 2022; Parisi et al., 2016), with subsequent extensions into offline learning contexts (Zhu et al., 2023; Lin et al., 2024) and methods to improve sample efficiency in these settings (Huang, 2022).
@@
Multi-policy MORL strategies directly target the approximation of the entire Pareto front by learning a diverse collection of policies. One direction for generating diverse behaviours involves developing single, highly adaptable models conditioned on preferences, which generalise across various objectives using techniques like specialised experience replay or policy gradient methods that enforce Pareto stationarity (Abels et al., 2019; Friedman and Fontaine, 2018; Kyriakis and Deshmukh, 2022). Other approaches explicitly learn a diverse set of policies or their value functions; this includes direct value-based methods like Pareto Q-learning (Van Moffaert and Nowé, 2014), and evolutionary algorithms often guided by prediction models to discover a dense Pareto set (Xu et al., 2020). Further techniques for generating policy sets involve Generalised Policy Improvement (GPI) for sample-efficient learning (Alegre et al., 2023) or the development of transferable policy components using representations like successor features (Alegre et al., 2022). The use of constrained optimisation to efficiently complete and refine the Pareto front is also explored in (Liu et al., 2024; He et al., 2024). Furthermore, the principles of decomposition-based strategies, which find a set of solutions by solving multiple interrelated scalarised sub-problems, have been a significant focus, with recent work providing clarifying taxonomies and conceptual frameworks (Felten et al., 2024; Röpke et al., 2024).
@@
While these established single-policy and multi-policy paradigms have significantly advanced MORL, the explicit characterisation and systematic exploitation of the structural relationship between the learned policies’ underlying parameter space and their resultant performance on the Pareto front remain largely underexplored. Although multi-objective optimization offers techniques for navigating Pareto sets (Ye and Liu, 2022), and some MORL studies have touched upon parameter space regularities (Xu et al., 2020), policy manifolds (Parisi et al., 2016), or front geometries (Li et al., 2024), these explorations typically do not formalise or exploit the parameter-to-performance mapping for systematic, guided Pareto front generation.
@@ 6 Conclusion
In this paper, we have introduced LLE-MORL, an algorithm that identifies solution components in multi-objective reinforcement learning. The main benefit of LLE-MORL is increased efficiency which is enabled by maintaining a direct relation between the multi-objective performance and the representation of the policy in the parameter space. Preference values of the objectives are used to generate a relatively small but diverse initial set of starting points in the weight space, but Pareto-optimality is not required for this initialisation, so that learning times can remain comparatively small. Likewise, brief retraining after linear shifts in the weight space compensates for non-linearities that are assumed to be small locally, so the current component of the Pareto front can be traced out efficiently. We have shown that this simple set-up is sufficient to obtain highly efficient coverage of a Pareto front which is superior to recent MORL algorithms. Although not shown here, it can be expected that the approach can be easily extended to more than $d=2$ objectives as the number of initial policy pairs (see Fig. 3) increases only linearly with $d$, although the representation of the policies set and of the Pareto front requires exponentially many points. An implicit representation of the Pareto front may seem to become preferable for $d>2$, although this would reduce the interpretability in terms of an accessible PPR as featured here.
Acknowledgments and Disclosure of Funding
@@ Appendix A Algorithm
In this section, we present a complete description of LLE-MORL, an efficient procedure for tracing an approximate Pareto front. Algorithm 1 details this process.
Algorithm 1 LLE-MORL
@@
1:Initial scalarization weights $\{w_{i}\}_{i=1}^{K}$ evenly spanning preference space, Target scalarization weights for directional retraining $\{w_{i}^{\prime}\}_{i=1}^{K}$, Initialization training length $T_{\mathrm{init}}$, Directional retraining length $T_{\mathrm{dir}}$, Fine-tuning length $T_{\mathrm{ref}}$, Step-scale factors $\{\alpha_{j}\}_{j=1}^{M}$
2:Approximate Pareto-optimal policy set $\Pi$
3:Initialization:
4:for $i=1$ to $K$ do
@@
5:     Train base policy $\theta_{w_{i}}$ with PPO under weight $w_{i}$ for $T_{\mathrm{init}}$ steps
6:end for
7:
8:Directional Retraining:
9:for $i=1$ to $K-1$ do
10:     $\theta_{w_{i}^{\prime}}\leftarrow$ continue training $\theta_{w_{i}}$ for $T_{\mathrm{dir}}$ steps under $w_{i}^{\prime}$
11:     $\Delta\theta_{i}\leftarrow\theta_{w_{i}^{\prime}}-\theta_{w_{i}}$
12:     $\Delta w_{i}\leftarrow w_{i}^{\prime}-w_{i}$
13:end for
14:
15:Locally Linear Extension:
16:$\mathcal{C}\leftarrow\emptyset$
17:for $i=1$ to $K-1$ do
18:     for $j=1$ to $M$ do
19:         $\theta_{i,j}\leftarrow\theta_{w_{i}}+\alpha_{j}\,\Delta\theta_{i}$
20:         $w_{i,j}\leftarrow w_{i}+\alpha_{j}\,\Delta w_{i}$
@@
21:         Evaluate performance $V(\theta_{i,j})$ under weight $w_{i,j}$
22:         $\mathcal{C}\leftarrow\mathcal{C}\cup\{(\theta_{i,j},w_{i,j})\}$
23:     end for
24:end for
25:
26:Candidate Selection:
27:$\mathcal{N}\leftarrow$ non-dominated subset of $\mathcal{C}$
28:
29:Preference-Aligned Fine-Tuning:
$\mathcal{F}\leftarrow\emptyset$
31:for all $(\theta,w)\in\mathcal{N}$ do

### v2

@@ 4.4 Ablation Study
The LLE-MORL integrates a locally linear extension process with a subsequent fine-tuning stage. To understand the distinct contributions of these components to the overall performance, our ablation study separates them. We first evaluate LLE-MORL-0, which solely employs the extension process without fine-tuning. As detailed in Table 1 and Table 2, LLE-MORL-0 itself demonstrates competitiveness, achieving strong HV and EU scores that are often competitive with or superior to baselines. This emphasises the efficacy of our extension mechanism in rapidly discovering a high-quality approximation of the Pareto front.
Subsequently, we assess the improvement of the fine-tuning stage by comparing LLE-MORL (which includes fine-tuning) to LLE-MORL-0. This comparison reveals that the inclusion of fine-tuning yields further improvements in HV and EU across all settings. The impact on SP is less uniform, which is an expected outcome, as refining solutions towards a more optimal Pareto front can alter their relative spacing. Nevertheless, the consistent enhancements in HV and EU prove the value of fine-tuning for improving the quality of the approximate Pareto front and its coverage by diverse solutions. This demonstrates that the extension process provides a strong foundation for the fine-tuning stage that enables LLE-MORL to outperform other algorithms.
@@ 5 Related work
Prior work in MORL offers various strategies for handling conflicting objectives. These can be broadly grouped into single-policy methods and multi-policy methods for approximating the Pareto front. Single-policy approaches typically convert the multi-objective problem into a single-objective task using a predefined preference or weighting scheme. A common instance of such a weighting scheme is linear scalarization  (Van Moffaert et al., 2013). Limitations of linear scalarization, particularly in capturing non-convex Pareto fronts, have been addressed by more advanced scalarization functions such as Chebyshev methods (Van Moffaert et al., 2013), hypervolume-based approaches (Zhang and Golovin, 2020) and the addition of concave terms to rewards (Lu et al., 2023). Concurrently, significant efforts have developed generalised single-policy models conditioned on preference inputs to achieve adaptability across diverse objectives (Teh et al., 2017; Yang et al., 2019; Basaklar et al., 2022; Parisi et al., 2016), with subsequent extensions into offline learning contexts (Zhu et al., 2023; Lin et al., 2024) and methods to improve sample efficiency in these settings (Hung et al., 2023).
@@
Multi-policy MORL strategies directly target the approximation of the entire Pareto front by learning a diverse collection of policies. One direction is developing single models conditioned on preferences (Abels et al., 2019; Friedman and Fontaine, 2018; Kyriakis and Deshmukh, 2022). Other approaches explicitly learn a diverse set of policies or their value functions; this includes direct value-based methods like Pareto Q-learning (Van Moffaert and Nowé, 2014), and evolutionary algorithms often guided by prediction models to discover a dense Pareto front (Xu et al., 2020). Further techniques for generating policy sets involve Generalised Policy Improvement (GPI) for sample-efficient learning (Alegre et al., 2023) or the development of transferable policy components using representations like successor features (Alegre et al., 2022). The use of constrained optimisation to efficiently complete and refine the Pareto front is also explored in (Liu et al., 2024; He et al., 2024). Furthermore, the principles of decomposition-based strategies, which find a set of solutions by solving multiple interrelated scalarised sub-problems, have been a significant focus (Felten et al., 2024; Röpke et al., 2024).
@@
While these single-policy and multi-policy paradigms have significantly advanced MORL, the characterisation and systematic exploitation of the structural relationship between the learned policies’ underlying parameter space and their resultant performance on the Pareto front remain largely underexplored. Although multi-objective optimisation offers techniques for Pareto navigation (Ye and Liu, 2022) and model merging (Rame et al., 2023; Chen and Kwok, 2025), and prior MORL studies have touched upon parameter space regularities (Xu et al., 2020), policy manifolds (Parisi et al., 2016; Shu et al., 2024), front geometries (Li et al., 2024), or preference control (Yuan et al., 2025; Yang et al., 2025), these directions do not provide a mechanism to systematically explore the local Pareto structure in a way that is interpretable and enables guided expansion.
@@ 6 Conclusion
We have discussed LLE-MORL, an algorithm that identifies solution components in multi-objective reinforcement learning by directed exploration of the Pareto front. The main benefit of LLE-MORL is interpretability and increased efficiency which is enabled by maintaining a direct relation between the multi-objective performance and the representation of the policy in the parameter space. We have shown that this simple set-up is sufficient to obtain a high-quality Pareto front from both empirical and theoretical aspects, and the LLE-MORL is superior to recent MORL algorithms. Furthermore, as the LLE stage is a modular plug-in component, it can be integrated into existing deep RL algorithms as a training-free and computationally lightweight module, offering a simple path for future extensions.
Acknowledgements
@@
Impact Statement
This paper aims to advance the field of multi-objective reinforcement learning by proposing a more efficient and interpretable algorithmic framework for approximating Pareto fronts. MORL could provide a flexible approach for adapting policies to user needs when these can be expressed through combinations of evaluative signals. We have shown that it is possible to achieve such an approach with strikingly less computational effort than the direct solutions, thereby reducing energy consumption and required computing resources. Likewise, improved results can be achieved in a shorter time, which can open the route to new applications e.g. in online process optimisation or control of complex systems, and enables the optimisation of more criteria in existing problems. The intended impact of this work is methodological, aimed at supporting research in reinforcement learning and multi-objective optimisation. As a general-purpose learning approach, the societal impacts of the proposed method depend on its application domain and on any side effects; we do not foresee any negative consequences within its intended scope.
@@ Appendix A Algorithm
Algorithm 1 LLE-MORL (General $d$-Objective Case)
@@
0: Initial scalarization weights $\{w_{k}\}_{k=1}^{K}$ evenly spanning the preference space, Directional retraining weights $\{w_{k}^{(i)}\}_{k=1,i=1}^{K,m}$, Initialization training length $T_{\mathrm{init}}$, Directional retraining length $T_{\mathrm{dir}}$, Fine-tuning length $T_{\mathrm{ref}}$, Step-scale factors $\alpha_{\mathrm{start}},\alpha_{\mathrm{end}},\Delta\alpha$
0: Approximate Pareto-optimal policy set $\Pi$
[…]
4:
5:Initialization:
6:for $k=1$ to $K$ do
@@
7:  Train base policy $\theta_{w_{k}}$ with PPO under weight $w_{k}$ for $T_{\mathrm{init}}$ steps
8:end for
9:
10:Directional Retraining:
11:for $k=1$ to K1K-1 do
12:  for $i=1$ to $m$ do
13:   $\theta_{k}^{(i)}\leftarrow$ continue training $\theta_{w_{k}}$ under $w_{k}^{(i)}$ for $T_{\mathrm{dir}}$ steps
14:   $\Delta\theta_{k}^{(i)}\leftarrow\theta_{k}^{(i)}-\theta_{w_{k}}$
15:   $\Delta w_{k}^{(i)}\leftarrow w_{k}^{(i)}-w_{k}$
16:  end for
17:end for
18:
19:Locally Linear Extension:
[…]
21:for $k=1$ to $K$ do
22:  for all combination of $\alpha_{i}$ on the grid $\alpha_{\mathrm{start}},\alpha_{\mathrm{end}},\Delta\alpha]$ do
23:   $\theta_{k,\mathrm{cand}}\leftarrow\theta_{w_{k}}+\sum_{i=1}^{m}\alpha_{i}\,\Delta\theta_{k}^{(i)}$
24:   $w_{k,\mathrm{cand}}\leftarrow w_{k}+\sum_{i=1}^{m}\alpha_{i}\,\Delta w_{k}^{(i)}$
@@
25:   Evaluate multi-objective performance $V(\theta_{k,\mathrm{cand}})$ under weight $w_{k,\mathrm{cand}}$
26:   $\mathcal{C}\leftarrow\mathcal{C}\cup\{(\theta_{k,\mathrm{cand}},w_{k,\mathrm{cand}})\}$
27:  end for
28:end for
29:
30:Candidate Selection:
[…]
32:
33:Preference-Aligned Fine-Tuning:
[…]
35:for all $(\theta,w)\in\mathcal{N}$ do